# Twin branching in shape memory alloys: a 1D continuum model with energy dissipation effects

### v1

@@ Appendix A Estimation of local interface propagation speed
The derivation is first performed for a simplified case when the (nominal) twin interfaces are perpendicular to the A–MM interface, thus ${\rm\bf m}\cdot{\rm\bf n}=0$ and $\alpha=1$, see Eq. (5)1. Consider thus a rigid translation of the unit cell with the speed $v_{\rm cell}$, as shown in Fig. 1(a). The resulting interface propagation speed $\hat{v}_{\rm n}$ (measured in the direction normal to the local twin interfaces) is thus given by
@@
Appendix B Computational treatment of the evolution problem
@@
The Lagrange functional $L$ is smooth (because function $l(\dot{h},\lambda)$ is continuously differentiable) hence the solution of the saddle-point problem (43) can be found by solving the condition of stationarity of $L$ with respect to arbitrary variations of $\dot{h}$ and $\lambda$, which reads
@@
Eq. (46) is a continuous counterpart of Eq. (32) with the effective Lagrange multiplier $\hat{\lambda}_{\rm eff}$ replacing the subdifferential $\partial_{\dot{h}}|\dot{h}|$. On the other hand, Eq. (47) is a weak form of a state-dependent constraint $C_{\rm eff}=0$ which enforces $\dot{h}=0$ when $|\hat{\lambda}|\leq 1$ and $\lambda=\operatorname{sign}\hat{\lambda}=\pm 1$ when $|\hat{\lambda}|>1$. Note that the augmented Lagrangian method delivers a numerically exact solution that is independent of the value of parameter $\varrho$.
@@
$\bar{\lambda}=\lambda+\bar{\varrho}\Delta\bar{h}$ and $\bar{\varrho}=L\varrho$. Note that the problem is rate-independent, hence, upon time discretization, the rate $\dot{h}$ is here approximated directly by the increment $\Delta h=h_{n+1}-h_{n}$ without specifying the corresponding time increment.

### v2

@@ Appendix B Estimation of local interface propagation speed
The derivation is first performed for a simplified case where the (nominal) twin interfaces are perpendicular to the A–MM interface, thus ${\rm\bf m}\cdot{\rm\bf n}=0$ and $\alpha=1$, see Eq. (5)1. Consider thus a rigid translation of the unit cell with the speed $v_{\rm cell}$, as shown in Fig. 1(a). The resulting interface propagation speed $\hat{v}_{\rm n}$ (measured in the direction normal to the local twin interfaces) is thus given by
@@
Appendix C Computational treatment of the evolution problem
@@
The Lagrange functional $L$ is smooth (because function $l(\dot{h},\lambda)$ is continuously differentiable) hence the solution of the saddle-point problem (53) can be found by solving the condition of stationarity of $L$ with respect to arbitrary variations $\delta\dot{h}$ and $\delta\lambda$, which reads
@@
Eq. (56) is a smooth counterpart of Eq. (34) with the effective Lagrange multiplier $\hat{\lambda}_{\rm eff}$ replacing the subdifferential $\partial_{\dot{h}}|\dot{h}|$. On the other hand, Eq. (57) is a weak form of a state-dependent constraint $C_{\rm eff}=0$ which enforces $\dot{h}=0$ when $|\hat{\lambda}|\leq 1$ and $\lambda=\operatorname{sign}\hat{\lambda}=\pm 1$ when $|\hat{\lambda}|>1$. Note that the augmented Lagrangian method delivers a numerically exact solution that is independent of the value of parameter $\varrho$.
@@
$\bar{\lambda}=\lambda+\bar{\varrho}\Delta\bar{h}$ and $\bar{\varrho}=L\varrho/\tau$. Note that the problem is rate-independent, hence, upon time discretization, the rate $\dot{h}$ is here approximated directly by the increment $\Delta h=h_{n+1}-h_{n}$ without specifying the corresponding time increment, and a constant $\tau=1$ s has been introduced to preserve the consistency of units.